実用技術英語 1 単位 (選択)

Practical Technical English  コインカー, パンカジ · マドゥカー · 特任助教 / 先端技術科学教育部

【授業目的】 This course is designed to help student to acquire the basic skills and

knowledge required for scientific writing. 【授業概要】 In this class student will a. learn strategies used in scientific writing

with an emphasis on effective argument, b. review student’s writing and revise their writing, c. give an oral presentation on student own technical paper

【キーワード】 Technical Writing 【先行科目】『上級技術英語』 (1.0) 【関連科目】『英語プレゼンテーション技法』 (1.0) 【履修要件】 You should have gotten the credit of ”Advanced Technical English.” 【履修上の注意】 This subject is a sompulsory one for Nichia STC students. If you

are not Nichia STC students and want to attend this class, please contact a coordinator in your department for getting permission of your attendance.

【到達目標】 1. To get ability to know the journalistic nuts and bolts of scientific writing

2. To get ability to discuss what science is, how to write a research paper,

3. To get ability to review topics of scientific importance

4. To get ability to distinguish central concepts from details in structuring scientific writing

5. To get ability to organize and present ideas logically

6. To get ability to write documents used in the sciences and engineering 【授業計画】

1. Basics of Scientific writing: Introduction to writing in the sciences and the challenges of scientific writing

2. Grammar and punctuation

3. The differences between scientific writing and other forms of writing

4. Describe several kinds of scientific writing and the purpose for each

5. Avoid common errors in scientific writing

6. Troublesome words and Commonly confused words

7. Getting to the point: Components of a research paper

8. Tips on writing Titles, Abstracts, and Introduction, Experimental Procedure

9. Tips on writing Results, Discussions, Conclusions, Acknowledgements and References

10. Visuals and their conventions 11. Writing for Publication: Grammar review of scientific research paper

12. Writing clearly and directly : Write precise and clear sentences and paragraphs 13. Improving scientific writing : Revise for clarity and effectiveness

14. Correcting grammatical flaws in sample texts

15. Ethics and scientific writing Fabrication, plagiarism, common knowledge, attributing authorship

16. Identify and correct common writing problems, Final revision questions and discussion

【成績評価基準】 The overall grade will be based on the mid-term and final examinations

and homework assignments. 【学習教育目標との関連】 [JABEE 関連] 【教科書】

⋄ S. Bailey, “Academic writing: A Practical guide for students” Nelson Thornes Ltd (2003) ⋄ Robert A. Day, Barbara Gastel, “How to write and publish a scientific paper”,

【参考書】 [参考資料]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215956

【対象学生】Nichia STC students and a limitted number of students who are

permitted to get this subject. 【連絡先】

⇒ コインカー (イノベーションセンター 3 階, 088-656-7643, koinkar@cicee.t

okushima-u.ac.jp) ���� ⇒ 橋爪 (E 棟 3 階南 D-2, 088-656-7473, tume@ee.tokushima-u.ac.jp) ����

Practical Technical English 1 unit (selection)

Pankaj Koinkar · Designated Assistant Professor / Graduate School of Advanced Technology and Science

Target⟩ This course is designed to help student to acquire the basic skills and

knowledge required for scientific writing. Outline⟩ In this class student will a. learn strategies used in scientific writing with

an emphasis on effective argument, b. review student’s writing and revise their writing, c. give an oral presentation on student own technical paper

�������⟩ Technical Writing Fundamental Lecture⟩ “Advanced Technical English”(1.0) Relational Lecture⟩ “Scientific Presentation Skills”(1.0) Requirement⟩ You should have gotten the credit of ”Advanced Technical English.” Notice⟩ This subject is a sompulsory one for Nichia STC students. If you are not

Nichia STC students and want to attend this class, please contact a coordinator in your department for getting permission of your attendance.

Goal⟩

1. To get ability to know the journalistic nuts and bolts of scientific writing

2. To get ability to discuss what science is, how to write a research paper,

3. To get ability to review topics of scientific importance

4. To get ability to distinguish central concepts from details in structuring scientific writing

5. To get ability to organize and present ideas logically

6. To get ability to write documents used in the sciences and engineering

Schedule⟩

1. Basics of Scientific writing: Introduction to writing in the sciences and the challenges of scientific writing

2. Grammar and punctuation

3. The differences between scientific writing and other forms of writing

4. Describe several kinds of scientific writing and the purpose for each

5. Avoid common errors in scientific writing

6. Troublesome words and Commonly confused words

7. Getting to the point: Components of a research paper

8. Tips on writing Titles, Abstracts, and Introduction, Experimental Procedure

9. Tips on writing Results, Discussions, Conclusions, Acknowledgements and References

10. Visuals and their conventions 11. Writing for Publication: Grammar review of scientific research paper

12. Writing clearly and directly : Write precise and clear sentences and paragraphs 13. Improving scientific writing : Revise for clarity and effectiveness

14. Correcting grammatical flaws in sample texts

15. Ethics and scientific writing Fabrication, plagiarism, common knowledge, attributing authorship

16. Identify and correct common writing problems, Final revision questions and discussion

Evaluation Criteria⟩ The overall grade will be based on the mid-term and final

examinations and homework assignments. Relation to Goal⟩ [JABEE 関連]

Textbook⟩

⋄ S. Bailey, “Academic writing: A Practical guide for students” Nelson Thornes Ltd (2003) ⋄ Robert A. Day, Barbara Gastel, “How to write and publish a scientific paper”,

Reference⟩ [参考資料]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215956

Student⟩ Nichia STC students and a limitted number of students who are permitted

to get this subject.

Contact⟩

⇒ Koinkar (イノベーションセンター 3 階, +81-88-656-7643, koinkar@cicee.t

okushima-u.ac.jp) ���� ⇒ Hashizume (E 棟 3 階南 D-2, +81-88-656-7473, tume@ee.tokushima-u.ac.j

p) ����